Find the impedance of the circuit if supply voltage is 

v(t) 110sin(20t 30 ) 220 cos(20t 90 )V

0 0

   

and the current 

i(t) 30cos(20t 60 ) 20sin(20t 60 )A


1
Expert's answer
2021-05-31T11:35:00-0400

Converting these sinusoids to the phasor domain to add them easily, 

"110sin (20t + 30°) + 220cos (20t — 90°) = 110\angle— 60° + 220\angle — 90°"

"= 55— j95.26— j220 = 55— j315.26 = 320.02\angle — 80.1°"  

Converting these sinusoids to the phasor domain to add them easily, 

"30cos (5t + 60°) — 20 sin (5t + 60°) = 30\angle — 30° — 20\angle 60°"

"= 15 + j25.98 — 17.321+ j10 = — 2.321+ j35.98 = 36.05\angle93.69°"

impedance


"Z=320/36\approx9 \ (V)"